SOLVE BY ELIMINATION BY ADDITION SHOW WORK PLEASE
7X-5Y=52
2X+5Y=47

Answer by fractalier(6550)

The idea is that we are allowed to add equations, and that's what we do to
7X-5Y=52
2X+5Y=47
If we add the, the 5y's drop out, they cancel, leaving
9X = 99 and
X = 11
Now we can plug that in to either equation and get
7(11) - 5Y = 52
-5Y = - 25
Y = 5